OpenStudy (anonymous):

Hello Considering positive values of x , locate the first two turning points. (a) y= 1-sinx (b) y=(1\2)x+ cos x I am concerned about the 2nd point on each equation / How to find it out ?

OpenStudy (anonymous):

i am not sure what "turning point" means here and why there should be two of them

OpenStudy (anonymous):

rather than in infinite number

OpenStudy (anonymous):

\[1-\sin(x)\] has a minimum value of \(0\) when \(\sin(x)=1\) i.e. when \(x=\frac{\pi}{2}\)

OpenStudy (anonymous):

oops lost @Lama97

OpenStudy (anonymous):

nop Iam here :) .. Turning points means stationary points :) well I got whats the min is but my question is asking for a max as well

OpenStudy (anonymous):

@satellite73 I hope you're still around :)

OpenStudy (anonymous):

yes the maximum value of \(1-\sin(x)\) is \(2\) when \(\sin(x)=-1\) i.e. when \(x=\frac{3\pi}{2}\)

OpenStudy (anonymous):

clear? not sure if this was supposed to be a calculus question but it certainly is not one the next may be though

OpenStudy (anonymous):

But when using differentiation and equating the differentiated eqn to zero I get only one point

OpenStudy (anonymous):

nop you're right :) its a calculus questin

OpenStudy (anonymous):

@satellite73 You're right my friend :)

OpenStudy (anonymous):

ok let me rant for a second there is no point in taking the derivative you get - cosine big deal you already know all about sine, so you do not need cosine but if you want to set cosine equal to zero and solve, go ahead, you still get \(\frac{\pi}{2}\) and \(\frac{3\pi}{2}\)

OpenStudy (anonymous):

how come I just got pie over two and not 3\2 pie :|

OpenStudy (anonymous):

cause you didn't look hard enough i guess cosine is zero two places on the interval \([0,2\pi)\)

OpenStudy (anonymous):

got it ":| Thanks for clearing it up

OpenStudy (anonymous):

this question required no calculus, but \[f(x)=\frac{1}{2}x+\cos(x)\] will

OpenStudy (anonymous):

derivative i think you got already as \[f'(x)=\frac{1}{2}-\sin(x)\] if you set it equal to zero you solve \[\sin(x)=\frac{1}{2}\] which occurs in more than on place in the interval \([0,2\pi)\)

OpenStudy (anonymous):

clear right? once at \(\frac{\pi}{6}\) and also at \(\frac{5\pi}{6}\)

OpenStudy (anonymous):

yeppp

OpenStudy (anonymous):

Thanks so much . Appreciate your help :)